OpenStudy (anonymous):

Equation : 2(12-8x)=x-11x

OpenStudy (fibonaccichick666):

Do you have any ideas on how to begin?

OpenStudy (anonymous):

Yes I do

OpenStudy (fibonaccichick666):

What do you think?

OpenStudy (anonymous):

I know how it's just the appear I have gives u the answer I don't know what I am doing wrongs

OpenStudy (fibonaccichick666):

ok well, will you walk through what you are doing and I'll tell you, please?

OpenStudy (fibonaccichick666):

Show every step no matter how small

OpenStudy (anonymous):

Ok I think that u would distribute first

OpenStudy (fibonaccichick666):

good and what do you get?

OpenStudy (anonymous):

24-16x=x-11x

OpenStudy (fibonaccichick666):

good, so next

OpenStudy (anonymous):

I think add 11x

OpenStudy (fibonaccichick666):

show me?

OpenStudy (anonymous):

16x+11x

OpenStudy (fibonaccichick666):

Stop, that's the spot

OpenStudy (fibonaccichick666):

you have : \[24-16x=x-11x\]\[24-16x+11x=x-11x+11x\] Can you simplify that?

OpenStudy (anonymous):

No????

OpenStudy (fibonaccichick666):

Where is the confusion?

OpenStudy (anonymous):

Ok so u can't simplify it u would have to subtract 11x from 16x

OpenStudy (fibonaccichick666):

you can simplify it, and you add because 11x is positive

OpenStudy (anonymous):

Ok that's what I thought then u would get 27x

OpenStudy (fibonaccichick666):

no, is it -16x or 16x

OpenStudy (anonymous):

Ok yeah then u would get 24=6x

OpenStudy (fibonaccichick666):

uhm... no

OpenStudy (fibonaccichick666):

ex. 4x+5x=9x 4x-5x=-x -4x+5x=x

OpenStudy (anonymous):

Yeah if you subtract. 16x from both sides instead of 11x

OpenStudy (fibonaccichick666):

no... you just do what I wrote

OpenStudy (fibonaccichick666):

\[24−16x+11x=x−11x+11x\] Simplify this using the above ex.

OpenStudy (anonymous):

24-27x=x

OpenStudy (fibonaccichick666):

how are you getting -27?